40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 28 19:08:50 2009, Yoichi, Update, PSL, Laser is back (sort of) Improving.png
    Reply  Thu Jan 29 13:52:34 2009, Yoichi, Update, PSL, Laser is back (sort of) SlowDC.pdf
Message ID: 1257     Entry time: Thu Jan 29 13:52:34 2009     In reply to: 1256
Author: Yoichi 
Type: Update 
Category: PSL 
Subject: Laser is back (sort of) 
Here is what I think has happened to the laser.

After the chiller line to the NPRO base clogged, the FSS slow slider went down to keep the laser frequency constant.
It is evident in the attachment 1 that the behavior of the slow slider and the DTEC (diode temp. stabilization feedback signal) are almost the same except for the direction. This means the slow servo was fighting against the increased heat caused by the lack of the cooling from the bottom.
DTEC was doing the same thing to keep the diode temperature constant.

Even though the slow actuator (a Peltier on the crystal) worked hard to keep the laser frequency constant, one can imagine that there was a large temperature gradient in the crystal and the mode shape may have changed.

Probably this made the coupling of the NPRO beam to the PA worse. It may also have put the NPRO in a mode hopping region, which could be the cause of the noisiness.

Right now, the MOPA power is 2.7W.
The FSS, PMC, MZ are locked. At first, the PMC locked on a sideband. I had to twiddle the phase flip button of the PMC servo to lock the PMC. Probably this is another sticky channel, which needs to be tweaked after a reboot of c1psl. I added a code to do this in /cvs/cds/caltech/scripts/Admin/slider_twiddle.

Currently the ISS is unstable. Kakeru and I are now taking OPLTF of the servo.
Looks like the phase margin at the lower UGF is too small.
Attachment 1: SlowDC.pdf  22 kB  | Hide | Hide all
SlowDC.pdf
ELOG V3.1.3-